Aluminum oxide can be formed from its elements.
4Al(s) + 3O2(g) → 2Al2O3(s)

What volume of oxygen is needed at 28.8 ∘C and 0.9450 atm  to completely react with 50.66 g of aluminum?

If 35.0 g of fluorine gas has a temperature of -20.3 ∘C and a pressure of 556 mmHg , what is its volume?

An airplane is pressurized to 650.0 mmHg , which is the atmospheric pressure at a ski resort at 1.30×104 ft altitude.

If air is 21.0 % oxygen, what is the partial pressure of oxygen on the plane?

If the partial pressure of oxygen drops below 100.0 mmHg , passengers become drowsy. If this happens, oxygen masks are released. What is the total cabin pressure at which oxygen masks are dropped?
Answer #1

4Al(s) + 3O2(g) → 2Al2O3(s)

4 moles of Al react 3 moles of O2

4*27g of Al react with 3*32g of O2

50.66g of Al react with = 3*32*50.66/4*27   = 45.03g of O2

PV = nRT

PV = WRT/M

V   = WRT/PM

      = 45.03*0.0821*301.8/0.945*32   = 36.9L of O2

2. P = 556mmHg   = 556/760   = 0.73atm

    T   = -20.3 +273 = 252.7K

   n = W/G.M.Wt

         = 35/38   = 0.92moles

PV = nRT

V = nRT/P

      = 0.92*0.0821*252.7/0.73   = 26.15L